OpenStudy (anonymous):

write an equation of the line whose slope is -2 and whose y-intercept is one

OpenStudy (ageta):

y = (m)x +(b) means y = (slope)x + (y intercept) y = (-1)x + (7) y = -x + 7

OpenStudy (ageta):

sorry that is not it

OpenStudy (anonymous):

y= mx+c m = slope c = y intercept so y = -2x+1

OpenStudy (dmndlife24):

y = mx + b where m = slope and b = y - intercept In the problem, m = -2, b = 1 y = -2x + 1

OpenStudy (ageta):

y = -2x + 1